| 12/16/2004 | ||||||||
| 1:30 | Still there is no data connection. | YS | ||||||
| 2:46 | data receaving started | TG | ||||||
| Bias is already on for CAL,HDS and SCD | ||||||||
| HV is at 3kV now. | ||||||||
| (HV of HDS's MB2 is 0. It looks strange.. Opher send email to Sonny about that) | ||||||||
| 3:29 | HV is ramped up to 6kV | TG | ||||||
| 3:45 | Bias of CAL MB5 is going up and down….. | TG | ||||||
| HV is still dead | ||||||||
| 4:03 | HV of CAL MB04 is going up and down at aroung 4.00kV | TG | ||||||
| It's temperature is higher than others (11.2) | ||||||||
| 4:08 | HV of CAL MB04 has been recovered. | TG | ||||||
| 4:11 | StandBy Mode | JH | ||||||
| 5:20 | data flow disconnected | TG | ||||||
| 5:45 | connected | JH | ||||||
| 5:56 | data receiving started | JH | ||||||
| 6:10 | disconnected | TG | <<< | Play-back | ||||
| 6:15 | connected (still in standby mode) | TG | <<< | at 2:40 on Dec 17 | ||||
| 7:01 | connection closed | KC | > there was no data files, no spi files. Only 600 events[0.2hz] | |||||
| 7:07 | data receiving started | KC | ||||||
| 7:50 | During rampping up HDS to 9kV MB08 is tripped | JH | ||||||
| 8:28 | MB5 in CAL has no trip (hv=3kv) | KC | ||||||
| 8:56 | BIOS=2.46 in MB5 (CAL) trip again(trip started about 8:30) | KC | ||||||
| 10:43 | Since14:28GMT we are in the flight mode due to log file record | KC | ||||||
| 13:27 | StandBy Mode | KC | ||||||
| 14:42 | Data disconnected-WFF trying re-establish | KC | <<< | play-back is done | ||||
| 14:54 | connected-data receiving rate increased to 86kbps | YS | <<< | at 2:14am on Dec 17. | ||||
| After reconnection, spi3 and pds are coming together now.. | ||||||||
| PKT0 is staying abound 14 -17% now | ||||||||
| 15:35 | Suddenly PKT0 went down to 0. (Maybe people on the ice sent command to clear buffer) | YS | ||||||
| pds files for CAL,HDS and SCD are not coming any more | ||||||||
| 16:04 | It became 'stand-by mode' by itself. | YS | ||||||
| We should run "daq_restart.stol" | ||||||||
| We are out of commanding window now… | ||||||||
| ESC said they are uploading parameters, we can't send commands now | ||||||||
| 16:20 | Sonny sent "stand-by mode" | YS | ||||||
| 17:04 | DATA flow stopped. No more housekeeping update. | <<< | Play-back is done with no data file. | |||||
| ESC said it'll be back in 5-10 minutes. | <<< | at 4:50am on DEC 17 | ||||||
| 17:18 | Housekeeping started updating | Maybe, it was stand-by mode.. | ||||||
| Connection comes back.. | ||||||||
| 17:57 | People on the ice uploading files… | |||||||
| ESC said they are goning to switch satellite. There will be a short lost connection. | ||||||||
| 18:10 | data folw reconnected | |||||||
| 18:36 | ICE uploading | |||||||
| 18:45 | LOS is terminated. All commanding will be thorough UMD only. | TG | ||||||
| 18:54 | ESC logined CDPS1 and made copy parameter files to their machine. | |||||||
| For this, we changed passwd temporarily and switch back to original | ||||||||
| 19:48 | Dr.Zinn sent some commands | |||||||
| 19.53 | Housekeeping continues to be updated, but data doesn`t flow anymore | |||||||
| 20:06 | We get data again | |||||||
| 20:30 | In case of stopping data taking and buffer clearing, follow this procedure. | TG | ||||||
| SCI_SW_OPMODE 0 | ||||||||
| SCI_SNIO_CONTROL 21,0 <==This command clears the buffer!! Please be careful |
||||||||
| Execute script daq_restart.stol | ||||||||
| 21:51 | we take over cmds. | TG | ||||||
| 22:14 | ERROR : master trigger : no data after 200001 attempts | |||||||
| going to standby mode | ||||||||
| According to Sonny`s instruction, we send following steps | ||||||||
| SCI_SW_OPMODE 0 | ||||||||
| SCI_SNIO_CONTROL 21,0 | ||||||||
| Execute script daq_restart.stol | ||||||||
| 23:15 | sent SCI_SW_NOOP | |||||||
| SCI_SW_OPMODE 0 | ||||||||
| SCI_SNIO_CONTROL 21,0 | ||||||||
| Execute script daq_restart.stol | ||||||||
| 23:28 | Warning : data file already open ; closing | |||||||
| And continue to download .pds files | ||||||||
| 23:40 | waited for commaning window to open | |||||||
| SCI_SW_OPMODE 0 | ||||||||
| SCI_SNIO_CONTROL 21,0 | ||||||||
| Execute script daq_restart.stol | ||||||||
| 23:46 | It went to standby mode automatically | |||||||
| We sent a cmd daq_restart.stol | ||||||||